OpenStudy (anonymous):

The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger?

OpenStudy (anonymous):

ok so 18 and older = x and 18 and younger = y 12x + 8y = 2,520 x + y = 224 right @pooja195

OpenStudy (igreen):

Correct!

OpenStudy (igreen):

Now, since 'y' is younger, solve for 'y'.

OpenStudy (anonymous):

ok give me a sec

OpenStudy (igreen):

We can re-arrange the 2nd equation: x + y = 224 Subtract 'y' to both sides: x = -y + 224 Plug in '-y + 224' for 'x' in the first equation: 12(-y + 224) + 8y = 2,520 Simplify and solve for 'y'

OpenStudy (anonymous):

Yes, Thats a good idea Green

OpenStudy (igreen):

Are you a teacher..?

OpenStudy (anonymous):

Yes I am.

OpenStudy (anonymous):

I teach 8th Grade Math

OpenStudy (igreen):

Oh, well you might want to read OpenStudy's \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/users/mlpsisters7}{Code~of~Conduct}}}}\).

OpenStudy (igreen):

I might be a Math Teacher when I grow up..I'm only 14 ATM though :P

OpenStudy (anonymous):

I teach 8th Grade Math

OpenStudy (anonymous):

I got x = 182 and y = 42 plug them in 12*182 + 8*42 = 2,520 2184 + 336 = 2520 2520 = 2520 182 + 42 = 224 224 = 224 there

sammixboo (sammixboo):

Right! =) Oh and \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\) @Mrs_Lopez

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Thank You very much @sammixboo

OpenStudy (igreen):

Yep, 42 is our answer then :P